OpenStudy (anonymous):

evaluate the function at the given values of x, if possible f(x)=6x^2-1/x^2 (a) x-=4 (b)x=-2

OpenStudy (anonymous):

is this \[f(x)=\frac{6x^2-1}{x^2}\]? if so then \[f(-2)=\frac{6\times (-2)^2-1}{(-2)^2}=\frac{24-1}{4}\]

OpenStudy (anonymous):

is a undefined